OpenStudy (anonymous):

The Kw of water varies with temperature. Calculate the pH of water at 30⁰C with a Kw = 1.469 x 10-14. Show all calculations leading to an answer. @girlover @iceicebaby

OpenStudy (anonymous):

plz i am so stuck

OpenStudy (anonymous):

Kw=[H+]*[OH-] Kw=x^2 x=[H+]=[OH-]=(1.469*(10^(-14)))^(1/2) pH=-log[H+]

OpenStudy (anonymous):

\[Kw=\sqrt{1.469\times10^{-14}}\]

OpenStudy (anonymous):

got it?

OpenStudy (cuanchi):

\[\left[ H ^{+} \right]=\sqrt{1.469 \times 10 ^{-14}}\] \[pH=-\log[H ^{+}]= - \log \sqrt{1.469 \times 10 ^{-14}}= 6.92\]

OpenStudy (anonymous):

i'm not a calculator but i think that manner is right.